OpenStudy (anonymous):

PLZ HELP I WILL MEDAL! Janine inflated 1 ball to a radius of 18 cm and another ball to a radius of 12 cm. How much more air was in the larger ball? Use 3.14 to approximate pi and express your answer in hundredths. ________cm3

OpenStudy (campbell_st):

the volume of a sphere is \[V = \frac{4}{3} \pi r^3\] calculate the volume of each ball r = 18 and r = 12 the find the difference

OpenStudy (campbell_st):

don't forget to use 3.14 in place of pi

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

oh one sec

OpenStudy (anonymous):

no i dont get it

OpenStudy (anonymous):

i really just need the answer bc school ends tomarrow at five and i have A CRAP TON of schoolwork to finish. and if i dont finish it i wont get into a military school i really want to go to

OpenStudy (welshfella):

air in larger balloon = 4/3 * 3.14 * 18^3 air in smaller balloon = 4/3 * 3.14 * 12^3

OpenStudy (campbell_st):

ok the large ball the volume will be \[V = \frac{4}{3} \times 3.14 \times 18^3\] the small ball \[V = \frac{4}{3} \times 3.14 \times 12^3\] do the 2 calculations and then find the difference...

OpenStudy (anonymous):

idk how to do it

OpenStudy (anonymous):

i got this so far -> 4/3* 18,312.4

OpenStudy (campbell_st):

do you have a calculator..?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

so if you got that far use \[4 \div 3 \times 18312.4\]

OpenStudy (anonymous):

24,416.5

OpenStudy (campbell_st):

that's good same thing for the small ball. then find the difference.... large ball - small ball

OpenStudy (anonymous):

i got 17,182 as the answer, is it correct?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

i have one more can you help

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

the answer was 17,182.08.